# Breaking the Reviewer: Assessing the Vulnerability of Large Language Models in Automated Peer Review Under Textual Adversarial Attacks

### v1

@@ 4.3 Evaluation – Review Quality
Table 1 shows the automatic evaluation results for the quality of reviews generated by LLMs. We begin by categorizing the human-written reviews into two types for comparison: (1) Within-Paper Pairs: review pairs from the same paper, representing the semantic similarity of reviews written for the same context. These pairs are generated using all possible combinations of reviews for each paper (e.g., $C^{n}_{2}$ combinations for $n$ reviews per paper), (2) Across-Paper Pairs: review pairs that are extracted from different papers, symbolizing the semantic similarity of reviews written for unrelated contents. We take all possible combinations of reviews across the entire dataset ($C^{n}_{2}$ combinations for N total human-written ground truth reviews) and exclude pairs that are included in the within-paper pairs to construct this set. Due to the significantly larger number of Across-Paper Pairs (2,774,190 pairs), we introduce random sampling to select 10% of the samples to create a subset (277,419 pairs), which is about 47 times larger than the 5,884 within-paper pairs. We observe that:
The ACOV of LLMs generated review is higher than human-written ground truths. The significance is further proved by Wilcoxon signed-rank test. The result can be attributed to the structured review guidelines provided in the prompt, which guide the model to cover a more comprehensive range of aspects. This suggests that LLMs have great potential to provide comprehensive perspectives on review principles in both positive and negative sentiments, demonstrating how LLMs can be guided to focus on desired qualities in reviews.
@@
Table 1: Evaluation of LLM-generated review’s Comprehensiveness and Similarity to human. ACOV stands for Aspect Coverage. R-1, R-2, R-L, and BS indicate the Rouge-1/2/L, and BertScore, respectively. Human (within) reports the mean scores obtained from all pairs of human reviews written for the same paper, while Human (across) reports the mean over pairs taken from different papers.
@@
Human (within) 0.5285 0.4148 0.0973 0.1899 0.8403
Human (across) 0.5285 0.2188 0.0285 0.1154 0.8468
GPT-4o-mini 0.6489 0.3938 0.820 0.1875 0.8341
GPT-4o 0.9151 0.3469 0.0678 0.1676 0.8251
Llama-3.3-70B 0.6489 0.3432 0.0750 0.1760 0.8208
Mistral-small-3.1 0.9202 0.3541 0.0737 0.1658 0.8220
Human-based review quality evaluation. While automatic metrics assess surface-level qualities like fluency and similarity to human-written text, they miss deeper aspects such as coherence, usefulness, and critical insight. To address this gap, we conduct a controlled human evaluation to complement the automatic analysis, as detailed in Appendix H.
@@ 4.4 Evaluation – Attack Effectiveness
Table 2: Adversarial Attack Performance and Score Shift Comparison of AgentReview Jin et al. (2024) dataset. “ASR” represents the Attack Success Rate. “Avg. Score shift” signifies the average score shift between attacked outputs and clean outputs. “Avg. #Pos/Neg Shift” indicates the average change in the number of positive and negative tags assigned to the reviews. “#Queries” refers to the average number of querying LLM. The value are reported as result from GPT-4o-mini / GPT-4o / Llama-3.3-70B / Mistral-Small-3.1.
Attacks ASR $\uparrow$ Average Shift #Queries
@@
DeepWordBug 0.87 / 0.90 / 0.81 / 0.91 3.44 / 4.41 / 2.56 / 3.55 0.24 / 0.01 / -0.08 / -0.02 0.78 / 1.49 / 1.70 / 0.32 43 / 51 / 52 / 55
PuncAttack 0.85 / 0.87 / 0.77 / 0.87 3.44 / 4.05 / 2.52 / 3.59 0.27 / 0.01 / -0.11 / -0.01 0.82 / 1.12 / 1.79 / 0.19 51 / 64 / 66 / 69
TextFooler 0.89 / 0.92 / 0.85 / 0.91 3.77 / 4.71 / 2.98 / 3.94 0.33 / 0.00 / -0.15 / -0.03 0.84 / 1.23 / 2.11 / 0.37 47 / 52 / 58 / 58
BERT-Attack 0.87 / 0.89 / 0.71 / 0.90 3.59 / 4.65 / 2.36 / 4.17 0.36 / -0.01 / -0.06 / -0.03 0.69 / 1.52 / 1.51 / 0.47 54 / 56 / 56 / 66
StyleAdv 0.89 / 0.92 / 0.82 / 0.92 3.62 / 4.73 / 2.87 / 3.97 0.30 / 0.01 / -0.16 / -0.04 0.82 / 1.31 / 2.25 / 0.38 25 / 32 / 25 / 30
As discussed in Section 3.3, we investigate the robustness of LLM-based reviewers by applying widely recognized textual adversarial attacks to manipulate input paper content. Table 2 summarizes the results of these attack baselines aimed at artificially inflating the review scores predicted by LLMs. We observe that:
Most attacks achieve high attack success rate (ASR). ASR among different granualarities of attacks ranging from 71% to 92%, accompanied by substantial score shifts. These findings demonstrate the effectiveness of adversarial methods in deceiving LLM-based reviewers and highlight the models’ sensitivity to such subtle manipulations injected into a lengthy document.
@@
Human-based attack imperceptibility evaluation. In addition to the quantitative results in Table 2, we conducted a user study to assess the stealthiness of attacks in a realistic peer review setting. Specifically, we evaluated participants’ ability to identify attacked papers and localize the perturbed paragraphs. The detailed settings and full user study results are reported in Appendix H.
5 Limitations of Possible Defenses
@@
Detecting Adversarial Manipulations in the Paper Content.
@@
Detecting Machine-generated Reviews.
@@
Policy Ambiguity. Due to current conference policies (e.g., the ACL Policy on Publication Ethics [^2] https://www.aclweb.org/adminwiki/index.php/ACL_Policy_on_Publication_Ethics#Guidelines_for_Generative_Assistance_in_Authorship and the NeurIPS 2025 LLM Policy [^3] https://neurips.cc/Conferences/2025/LLM) explicitly allowing the use of LLMs for language refinement, there is no clear boundary between LLM-written and LLM-assisted reviews. Applying a strict detection threshold risks unfairly penalizing legitimate human-written reviews that simply use LLMs for improving wording and fluency.
Ease of Bypassing Detection. Existing detectors are vulnerable under various real-world application scenarios. An adversary can paraphrase LLM-generated reviews Krishna et al. (2023) or design prompts to guide LLMs in generating human-like text to bypass detectors Wu et al. (2024), while these reviews still rely on LLM content that lacks a deep understanding of the paper. This undermines the goal of ensuring high-quality reviews. Futhermore, as conferences begin to officially adopt LLM tools in the review pipeline (e.g., AAAI’s AI-powered peer review assessment system [^4] https://aaai.org/aaai-launches-ai-powered-peer-review-assessment-system/), relying solely on detector-based filtering appears inadequate to address the vulnerabilities introduced by adversarial manipulations of the paper content.
@@ 6 Conclusion
We introduce a framework for evaluating the capabilities and robustness of large language models (LLMs) in automated peer-review tasks, addressing the significant burden on reviewers from the surge in manuscript submissions. Utilizing mainstream LLMs, we perform two core tasks—review generation and score prediction—demonstrating the effectiveness of LLMs in providing coherent and insightful feedback. To assess robustness, we apply various adversarial techniques to manipulate paper content and observe their impact on the LLM’s predicted scores, revealing vulnerabilities where subtle input changes significantly affect reliability. In response, we plan to develop defense mechanisms, such as inference-time detection methods like perplexity detection, to enhance the stability of LLM-based peer review processes.
@@ 8 Ethical Considerations
Fragility of AI-text Detectors.  While utilizing AI-text detectors to identify machine-generated content appears to be a potential defense against our proposed threats, such mechanisms are insufficient to ensure security in real-world peer reviewing scenario for several reasons (detailed in Section 5).
First, current detectors are primarily trained to distinguish between human-written and LLM-generated text, but they lack the capability to detect subtle, semantics-preserving adversarial perturbations in paper content. Second, zero-shot detection of LLM-generated reviews can be trivially bypassed through paraphrasing or minor editing, allowing AI-text to evade filtering. Moereover, the increasing integration of LLMs into legitimate writing and reviewing workflows—often under ambiguous policy guidelines—renders binary detections ethically problematic and operationally difficult.
@@
Appendix A Prompt Example
Fig. 4 presents the detailed prompt template used for automatic review generation in our experiments. The prompt is explicitly crafted to simulate the role of a professional reviewer specialized in computer science and machine learning, guiding LLMs to produce structured reviews in ICLR style. To ensure comprehensive critique, the prompt directs the model to include both positive and negative aspects of the paper by emphasizing: “Make sure to include both positive and negative aspects.” This approach addresses findings from Zhou et al. (2024), who noted that LLMs often generate overly positive reviews, risking optimistic evaluations that overlook a paper’s weaknesses. By encouraging balanced feedback, the prompt aims to produce more critical and informative assessments.
Prompt Template You are a professional reviewer in computer science and machine learning. Based on the given content of a research paper, you need to write a review in ICLR style and tag sentences with the corresponding tag type at the beginning of sequence: tags types: [NONE], [SUMMARY], [MOTIVATION POSITIVE], [MOTIVATION NEGATIVE], [SUBSTANCE POSITIVE], [SUBSTANCE NEGATIVE], [ORIGINALITY POSITIVE], [ORIGINALITY NEGATIVE], [SOUNDNESS POSITIVE], [SOUNDNESS NEGATIVE], [CLARITY POSITIVE], [CLARITY NEGATIVE], [REPLICABILITY POSITIVE], [REPLICABILITY NEGATIVE], [MEANINGFUL COMPARISON POSITIVE], [MEANINGFUL COMPARISON NEGATIVE]. Your total output should not surpass 500 tokens, make sure to include both positive and negative aspects. Also, you need to predict the review score in several aspects based on the generated review, providing an explanation of each aspect in less than 30 tokens. Choose a integer score from 1 to 10, higher score means better paper quality. Section {section name 1}: {section content 1} Section {section name 2}: {section content 2} Section {section name N}: {section content N} Please strictly follow the format of Example output: 1. REVIEW: tagged sequences. 2. REVIEW SCORE: OVERALL: score, SUBSTANCE: score, APPROPRIATENESS: score, MEANINGFUL COMPARISON: score, SOUNDNESS CORRECTNESS: score, ORIGINALITY: score, CLARITY: score, IMPACT: score. 3. REVIEW SCORE EXPLANATION: OVERALL: explanation, SUBSTANCE: explanation, APPROPRIATENESS: explanation, MEANINGFUL COMPARISON: explanation, SOUNDNESS CORRECTNESS: explanation, ORIGINALITY: explanation, CLARITY: explanation, IMPACT: explanation.
Figure 4: Prompt template used for review generation.
Appendix B Implementation Details – Model Configuration
We accessed closed-source models (GPT-4o and GPT-4o-mini) via the OpenAI API, and open-source models—Llama-3.3-70B (approximately 70 billion parameters, running on H100 via vLLM, requiring around 140GB of memory) and Mistral-Small-3.1 (approximately 24 billion parameters, running on H100 via vLLM, requiring around 60GB)—through Hugging Face. All models were configured with a temperature of 0.3 to balance creativity and coherence. The maximum number of generated tokens was set to 2048, ensuring detailed reviews without exceeding practical length constraints.
Appendix C Case Study
@@
Table 3: Aspect score predicted by GPT-4o-mini with the explanations.
@@
Meaningful Comparison 7.0 The paper contextualizes its contributions well but could engage more critically with alternative approaches.
Soundness Correctness 8.0 The theoretical framework is robust, though some assumptions may limit applicability.
@@
Appendix D Implementation Details – Longgest Common Sequence (LCS) split
@@
Character-level Attack: The LCS algorithm then identifies the longest common subsequences at the character fine-grained level. This allows for precise modifications at the character level, which is useful for attacks that require small, detailed changes.
Word-level Attack: For a word-level attack, both the clean paper content and the generated review are split by individual words. The LCS matching process is constrained to operate on these word-level splits, as word-level attacks typically involve replacing words with synonyms or other word-based modifications. This approach ensures that each modifiable subsequence corresponds to complete words, enabling word-level changes to occur seamlessly.
Sentence-level Attack: For sentence-level attacks, the LCS is first performed at the word level, identifying the longest common subsequences between words in the clean content and the review. Once the LCS matches are found at the word level, the matched sequence is expanded to include the full sentence. Specifically, for any word sequence that forms a matched subsequence, we extend the match forward and backward to capture the entire sentence in which the sequence appears. This ensures that the adversarial patterns can be injected into complete sentences, preserving the overall syntactic and semantic structure of the content.
Appendix E Implementation Details – Textual Adversarial Attack.
Textual attacks are composed of four fundamental components: (1) Goal Function defines the objective of the attack, offering a heuristic for search strategies to identify optimal solutions. It evaluates how effectively perturbations achieve desired outcomes, such as convincing LLMs to favor the altered paper for better ratings and feedback. (2) Constraints ensure that the disturbed text maintains a resemblance to the original, preserving semantic similarity to validate the modification. (3) Transformation process details methods for altering input content, including character modifications, synonym replacements, and changes in style and syntax. (4) Search Method explores valid perturbations produced by transformations, employing techniques such as greedy search with word importance ranking, beam search, and brute-force search.
For the search part, we adopt a greedy approach with word importance ranking for DeepWordBug, BERT-Attack, PuncAttack, and TextFooler. We restrict our process to the top 50 most important words to balance computational affordability and attack effectiveness, feeding these into our perturbation generation pipeline. StyleAdv uses brute-force at the sentence level because both methods do not involve in word-level searching. We segment paper content into individual sentences using the NLTK tokenizer [^6] https://www.nltk.org/api/nltk.tokenize.html Bird (2006). For style attacks, we utilize the pre-trained model Krishna et al. (2020) to convert text into Bible style. Original sentences are replaced with their transformed counterparts as adversarial examples.
Appendix F Degree of changes using different adversarial attacks
To further quantify the degree of changes when applying different textual adversarial attacks on paper content, we report two standard metrics Modification Rate and Semantic Similarity.
@@ Appendix F Degree of changes using different adversarial attacks
$\mathcal{S}=\big{\{}(s_{\text{clean}},s_{\text{adv}})\ \big{|}$ $s_{\text{clean}}\in x_{\text{clean}},s_{\text{adv}}\in x_{\text{adv}},$
@@
Each pair $(s_{\text{clean}},s_{\text{adv}})\in\mathcal{S}$ contains a sentence from the original and its corresponding modified version from the adversarial input. We then compute two metrics over this set.
  • Modification Rate, which measures the extent of perturbation introduced into the original text.
    For each sentence pair in $(s_{\text{clean}},s_{\text{adv}})\in\mathcal{S}$ , the normalized character-level change is:
    Modification $\text{Modification Rate}=\frac{\Delta_{\text{char}}(s_{\text{clean}},s_{\text{adv}})}{(|s_{\text{clean}}|+|s_{\text{adv}}|)/2}$
    where $\Delta_{\text{char}}(s_{\text{clean}},s_{\text{adv}})$ denotes the number of changed characters (insertions, deletions, or substitutions).
  • Semantic Similarity, measured by BERTScore, to assess how closely the adversarial texts $s_{\text{adv}}$ preserve the original semantics $s_{\text{clean}}$
@@
In addition, to qualitatively evaluate the degree of changes, we have provided more examples in Table 5 shown below.
From these examples in the table, we can find that character-level attacks like DeepWordBug and PuncAttack make minimal, often imperceptible changes that nonetheless disrupt model predictions, revealing vulnerability to small spelling errors. Word-level attacks (TextFooler, BERT-Attack) replace words with adversarial synonyms, maintaining syntactic structure and semantics but misleading models reliant on lexical cues. Sentence-level attacks (StyleAdv) introduce broader stylistic and phrasing changes, preserving overall meaning but posing challenges for detection. These results highlight a trade-off between semantic preservation and attack stealthiness, emphasizing the need for defenses that handle perturbations from character edits to stylistic rewrites.
@@
Appendix G Evaluation – Adversarial Attack Result for PeerRead Dataset

### v3

@@ 4.3 Evaluation – Review Quality
Table 1 reports automatic evaluation results for the quality of reviews generated by LLMs. To establish comparison baselines, we first categorize human-written reviews into two types: (1) Within-Paper Pairs: review pairs originating from the same paper, capturing the semantic similarity of reviews written for the same context. These pairs are generated using all possible combinations of reviews for each paper (e.g., $C^{n}_{2}$ combinations for $n$ reviews per paper). (2) Across-Paper Pairs: review pairs formed from different papers, reflecting the semantic similarity of reviews written for unrelated content. This set is constructed by enumerating all possible combinations of reviews across the entire dataset ($C^{n}_{2}$ combinations for N total human-written reviews) and excluding those counted as within-paper pairs. [^2] Since Across-Paper Pairs are substantially more numerous, we randomly sample 10% for similarity calculation to reduce computational cost. Unlike similarity-based metrics, ACOV does not rely on paired data, we do not split it into within- and across-paper categories. Consequently, its score for human-written reviews is identical across pair types. We observe that:
The ACOV of LLM-generated reviews is higher than human-written ground truths. The significance is further proved by the Wilcoxon signed-rank test. The result can be attributed to the structured review guidelines provided in the prompt, which guide the model in covering a more comprehensive range of aspects. These findings suggest that, under explicit guidance, LLMs can provide broader perspectives on review principles, covering both positive and negative sentiments while focusing on the desired qualities of reviews. Importantly, our intention is not to claim that LLM-generated reviews surpass human reviews, but rather to highlight their potential for offering evaluations across multiple perspectives. We also note that the ACOV decreases when the scoring aspects are not explicitly specified, due to the lack of direct alignment. A detailed interpretation is provided in Appendix I.
@@
Table 1: Evaluation of LLM-generated reviews’ Comprehensiveness and Similarity to humans of AgentReview Jin et al. (2024) dataset. ACOV denotes Aspect Coverage. R-1, R-2, R-L, and BS indicate the Rouge-1/2/L, and BertScore, respectively. Human (within) reports the mean scores of within-paper pairs, Human (across) reports the mean over across-paper pairs.
@@
Human (within) 0.4980 0.3746 0.0774 0.1813 0.8440
Human (across) 0.4980 0.1994 0.0292 0.1069 0.8163
GPT-4o-mini 0.6111 0.3185 0.0670 0.1594 0.8285
GPT-4o 0.9162 0.3134 0.0666 0.1574 0.8254
Llama-3.3-70B 0.6594 0.3127 0.0750 0.1666 0.8272
Mistral-small-3.1 0.9149 0.3204 0.0731 0.1552 0.8226
@@ 4.4 Evaluation – Attack Effectiveness
Table 2: Adversarial Attack Performance and Score Shift Comparison of AgentReview Jin et al. (2024) dataset. “ASR” represents the Attack Success Rate. “Avg. Score shift” signifies the average score shift between attacked and clean outputs. “Avg. #Pos/Neg Shift” indicates the average change in the number of positive and negative tags assigned to the reviews. “#Queries” refers to the average number of times LLM is queried. The values are reported as result from GPT-4o-mini / GPT-4o / Llama-3.3-70B / Mistral-small-3.1.
Attacks ASR $\uparrow$ Average Shift #Queries $\downarrow$
@@
DeepWordBug 0.87 / 0.90 / 0.81 / 0.91 3.44 / 4.41 / 2.56 / 3.55 0.24 / 0.01 / -0.08 / -0.02 -0.78 / -1.49 / -1.70 / -0.32 43 / 51 / 52 / 55
PuncAttack 0.85 / 0.87 / 0.77 / 0.87 3.44 / 4.05 / 2.52 / 3.59 0.27 / 0.01 / -0.11 / -0.01 -0.82 / -1.12 / -1.79 / -0.19 51 / 64 / 66 / 69
TextFooler 0.89 / 0.92 / 0.85 / 0.91 3.77 / 4.71 / 2.98 / 3.94 0.33 / 0.00 / -0.15 / -0.03 -0.84 / -1.23 / -2.11 / -0.37 47 / 52 / 58 / 58
BERT-Attack 0.87 / 0.89 / 0.71 / 0.90 3.59 / 4.65 / 2.36 / 4.17 0.36 / -0.01 / -0.06 / -0.03 -0.69 / -1.52 / -1.51 / -0.47 54 / 56 / 56 / 66
StyleAdv 0.89 / 0.92 / 0.82 / 0.92 3.62 / 4.73 / 2.87 / 3.97 0.30 / 0.01 / -0.15 / -0.04 -0.82 / -1.31 / -2.25 / -0.38 25 / 32 / 25 / 30
To assess the robustness of LLM-based reviewers, we employ widely recognized textual adversarial attacks (described in Section 3.3) to manipulate the content of input paper. Table 2 summarizes the results of these attack baselines aimed at artificially inflating the review scores predicted by LLMs. We observe that:
Most attacks achieve a high attack success rate (ASR). ASR among different granularities of attacks ranging from 71% to 92%, accompanied by substantial score shifts. These findings demonstrate the effectiveness of adversarial methods in deceiving LLM-based reviewers and highlight the models’ sensitivity to such subtle manipulations injected into a lengthy document.
@@
5 Defenses and Limitations
@@
5.1 Detecting Adversarial Manipulations.
@@
5.2 Detecting Machine-Generated Reviews.
@@
Policy Ambiguity. Due to current conference policies (e.g., the ACL Policy on Publication Ethics [^3] https://www.aclweb.org/adminwiki/index.php/ACL_Policy_on_Publication_Ethics#Guidelines_for_Generative_Assistance_in_Authorship and the NeurIPS 2025 LLM Policy [^4] https://neurips.cc/Conferences/2025/LLM) explicitly allowing the use of LLMs for language refinement, there is no clear boundary between LLM-written and LLM-assisted reviews. Applying a strict detection threshold risks unfairly penalizing legitimate human-written reviews that simply use LLMs to improve wording and fluency.
Ease of Bypassing Detection. Existing detectors are vulnerable under various real-world application scenarios. An adversary can paraphrase LLM-generated reviews Krishna et al. (2023) or design prompts to guide LLMs in generating human-like text to bypass detectors Wu et al. (2024), while these reviews still rely on LLM content that lacks a deep understanding of the paper. This undermines the goal of ensuring high-quality reviews. Furthermore, as conferences begin to officially adopt LLM tools in the review pipeline (e.g., AAAI’s AI-powered peer review assessment system [^5] https://aaai.org/aaai-launches-ai-powered-peer-review-assessment-system/), relying solely on detector-based filtering appears inadequate to address the vulnerabilities introduced by adversarial manipulations of the paper content.
@@ 6 Conclusion
We present a framework to evaluate the capabilities and robustness of large language models (LLMs) in automated peer-review tasks, addressing the growing burden on reviewers. Using mainstream LLMs, we perform review generation and score prediction, demonstrating their ability to provide coherent, insightful feedback. To assess robustness, we apply adversarial techniques to manipulate paper content and observe their impact on predicted scores, revealing vulnerabilities where subtle input changes affect reliability. We also propose Attack Focus Localization to identify document regions most vulnerable to attacks, offering finer-grained insights into LLM weaknesses. Although such manipulations remain difficult to detect, exploring mitigation strategies—such as inference-time monitoring or human-in-the-loop verification—represents a promising avenue to improve the trustworthiness of LLM-assisted peer review.
Acknowledgments
This work was supported in part by the National Science and Technology Council (NSTC) of Taiwan under Grant NSTC-112-2221-E-A49-094-MY3, and in part by the U.S. National Science Foundation (NSF) under Grants III-2106758 and POSE-2346158.
@@ 8 Ethical Considerations
Fragility of AI-text Detectors.  While utilizing AI-text detectors to identify machine-generated content appears to be a potential defense against our proposed threats, such mechanisms are insufficient to ensure security in real-world peer reviewing scenarios for several reasons (detailed in Section 5).
First, current detectors are primarily trained to distinguish between human-written and LLM-generated text, but they lack the capability to detect subtle, semantics-preserving adversarial perturbations in paper content. Second, zero-shot detection of LLM-generated reviews can be trivially bypassed through paraphrasing or minor editing, allowing AI-text to evade filtering. Moreover, the increasing integration of LLMs into legitimate writing and reviewing workflows—often under ambiguous policy guidelines—renders binary detections ethically problematic and operationally difficult.
@@
Appendix A Case Study
@@
Table 3: Predicted aspect scores from GPT-4o-mini and their corresponding explanations.
@@
Meaningful
Comparison
7.0 The paper contextualizes its contributions well but could engage more critically with alternative approaches.
Soundness
Correctness
8.0 The theoretical framework is robust, though some assumptions may limit applicability.
@@
Appendix B Granularity-Specific LCS Splitting
@@
Character-level Splitting. The LCS algorithm operates at a fine-grained character level, enabling precise modifications that are particularly useful for attacks requiring small, detailed changes.
Word-level Splitting. We split both the paper content and the generated review into word units and constrain the LCS matching to operate on those units. Since word-level attacks typically involve synonym substitutions or similar changes in word granularity, this ensures modifiable subsequences correspond to complete words and that replacements do not fragment tokens.
Sentence-level Splitting. For sentence-level attacks, we first apply LCS at the word level to find matched subsequences between the clean content and the review, then extend each match bidirectionally to capture the full sentence containing the sequence. This ensures adversarial patterns are injected into complete sentences while preserving the overall syntactic and semantic coherence.
Appendix C Magnitude of Changes Across Different Adversarial Attacks
To quantify the magnitude of changes when applying different textual adversarial attacks on paper content, we report two standard metrics: Modification Rate and Semantic Similarity.
@@ Appendix C Magnitude of Changes Across Different Adversarial Attacks
$\mathcal{S}=\big\{(s_{\text{clean}},s_{\text{adv}})\ \big|\$$s_{\text{clean}}\in x_{\text{clean}},s_{\text{adv}}\in x_{\text{adv}},$
@@
Each pair in $\mathcal{S}$ consists of an original sentence and its corresponding adversarial variant. The following two metrics are computed over this set:
Modification Rate. This metric measures the extent of perturbation introduced into the original text. For each sentence pair in $\mathcal{S}$, the normalized character-level change is:
$\text{Modification Rate}=\frac{\Delta_{\text{char}}(s_{\text{clean}},s_{\text{adv}})}{(|s_{\text{clean}}|+|s_{\text{adv}}|)/2}$
where $\Delta_{\text{char}}(s_{\text{clean}},s_{\text{adv}})$ denotes the total number of characters that differ between $s_{\text{clean}}$ and $s_{\text{adv}}$, counting insertions, deletions, and substitutions.
Semantic Similarity. This metric, measured by BERTScore, assesses how closely the adversarial text $s_{\text{adv}}$ preserves the semantics of $s_{\text{clean}}$.
@@
In addition, we provide examples in Table 5 to qualitatively evaluate the degree of changes. These examples show that character-level attacks (DeepWordBug, PuncAttack) make minimal, often imperceptible changes that nonetheless disrupt model predictions, revealing vulnerability to small spelling errors. Word-level attacks (TextFooler, BERT-Attack) replace words with adversarial synonyms, maintaining syntactic structure and semantics but misleading models reliant on lexical cues. Sentence-level attacks (StyleAdv) introduce broader stylistic and phrasing changes, preserving overall meaning while posing challenges for detection. These results highlight how different adversarial attacks vary in modification magnitude and detectability, emphasizing the need for defenses that handle perturbations from minor character edits to broader stylistic rewrites.
@@
Appendix D Components and Implementation of Textual Adversarial Attacks.
Textual attacks are composed of four fundamental components: (1) Goal Function defines the objective of the attack, offering a heuristic for search strategies to identify optimal solutions. It evaluates how effectively perturbations achieve desired outcomes, such as convincing LLMs to favor the altered paper for better ratings and feedback. (2) Constraints ensure that the disturbed text maintains a resemblance to the original, preserving semantic similarity to validate the modification. (3) Transformation process details methods for altering input content, including character modifications, synonym replacements, and changes in style. (4) Search Method explores valid perturbations produced by transformations, employing techniques such as greedy search with word importance ranking, beam search, and brute-force search.
For the search part, we adopt a greedy approach with word importance ranking for DeepWordBug, BERT-Attack, PuncAttack, and TextFooler. We restrict our process to the top 50 most important words to balance computational affordability and attack effectiveness, feeding these into our perturbation generation pipeline. StyleAdv uses brute-force at the sentence level because it does not involve word-level searching. We segment paper content into individual sentences using the NLTK tokenizer [^7] https://www.nltk.org/api/nltk.tokenize.html Bird (2006). For style attacks, we utilize the pre-trained model Krishna et al. (2020) to convert text into Bible style. Original sentences are replaced with their transformed counterparts as adversarial examples.
Appendix E Model Configuration
We accessed closed-source models (GPT-4o and GPT-4o-mini) via the OpenAI API. Open-source models were accessed through Hugging Face, including Llama-3.3-70B (approximately 70B parameters, running on H100 via vLLM, requiring around 140GB memory) and Mistral-small-3.1 (approximately 24B parameters, running on H100 via vLLM, requiring around 60GB memory). All models were configured with a temperature of 0.3 to balance creativity and coherence. The maximum number of generated tokens was set to 2048, allowing detailed reviews without exceeding practical length constraints.
Appendix F Prompt Example
Fig. 4 presents the detailed prompt template used for automatic review generation in our experiments. The prompt is explicitly crafted to simulate the role of a professional reviewer specialized in computer science and machine learning, guiding LLMs to produce structured reviews in ICLR style. To ensure a comprehensive critique, the prompt directs the model to include both positive and negative aspects of the paper by emphasizing: “Make sure to include both positive and negative aspects.” This approach addresses findings from Zhou et al. (2024), who noted that LLMs often generate overly positive reviews, risking optimistic evaluations that overlook a paper’s weaknesses. By encouraging balanced feedback, the prompt aims to produce more critical and informative assessments.
Prompt Template You are a professional reviewer in computer science and machine learning. Based on the given content of a research paper, you need to write a review in ICLR style and tag sentences with the corresponding tag type at the beginning of sequence: tags types: [NONE], [SUMMARY], [MOTIVATION POSITIVE], [MOTIVATION NEGATIVE], [SUBSTANCE POSITIVE], [SUBSTANCE NEGATIVE], [ORIGINALITY POSITIVE], [ORIGINALITY NEGATIVE], [SOUNDNESS POSITIVE], [SOUNDNESS NEGATIVE], [CLARITY POSITIVE], [CLARITY NEGATIVE], [REPLICABILITY POSITIVE], [REPLICABILITY NEGATIVE], [MEANINGFUL COMPARISON POSITIVE], [MEANINGFUL COMPARISON NEGATIVE]. Your total output should not surpass 500 tokens, make sure to include both positive and negative aspects. Also, you need to predict the review score in several aspects based on the generated review, providing an explanation of each aspect in less than 30 tokens. Choose an integer score from 1 to 10, higher score means better paper quality. Section {section name 1}: {section content 1} Section {section name 2}: {section content 2} Section {section name N}: {section content N} Please strictly follow the format of Example output: 1. REVIEW: tagged sequences. 2. REVIEW SCORE: OVERALL: score, SUBSTANCE: score, APPROPRIATENESS: score, MEANINGFUL COMPARISON: score, SOUNDNESS CORRECTNESS: score, ORIGINALITY: score, CLARITY: score, IMPACT: score. 3. REVIEW SCORE EXPLANATION: OVERALL: explanation, SUBSTANCE: explanation, APPROPRIATENESS: explanation, MEANINGFUL COMPARISON: explanation, SOUNDNESS CORRECTNESS: explanation, ORIGINALITY: explanation, CLARITY: explanation, IMPACT: explanation.
Figure 4: Prompt template used for review generation.
Appendix G Experiment Result of PeerRead Dataset
G.1 Evaluation – Review Quality
The evaluation of review quality of PeerRead dataset is provided in Table 6, which exhibits trends consistent with those discussed in Section 4.3.
Table 6: Evaluation of LLM-generated reviews’ Comprehensiveness and Similarity to humans of PeerRead Kang et al. (2018) dataset. ACOV denotes Aspect Coverage. R-1, R-2, R-L, and BS indicate the Rouge-1/2/L, and BertScore, respectively. Human (within) reports the mean scores of within-paper pairs, Human (across) reports the mean over across-paper pairs.
Reviewer ACOV R-1 R-2 R-L BS
Human (within) 0.5285 0.4148 0.0973 0.1899 0.8403
Human (across) 0.5285 0.2188 0.0285 0.1154 0.8179
GPT-4o-mini 0.6066 0.3938 0.0676 0.1875 0.8341
GPT-4o 0.9151 0.3469 0.0678 0.1676 0.8251
Llama-3.3-70B 0.6489 0.3432 0.0757 0.1760 0.8208
Mistral-small-3.1 0.9202 0.3541 0.0737 0.1658 0.8220
G.2 Evaluation – Attack Effectiveness